what would x be is the surface area is 864 cm2 and height is 12 cm and length is 16 cm
Answer by ewatrrr(24785)

 
Hi,

Surface = 2(12*16) + 2(12x) + 2(21x) = 664cm^2
2(12x) + 2(21x) = 664-384
66x = 280
x = 4.24